A page containing the append function. [append →^{pyk} “append”] [append(x, y) →^{tex} “append(#1., #2.)”] [append(x, y) →^{val} y : : x] [append(2, 1) ≈ 1 : : 2]^{·} [append(3, 1 : : 2) ≈ 1 : : 2 : : 3]^{·}